# Bayesian Speech synthesizers Can Learn from Multiple Teachers

### v1

@@ C.1 Training Data
To provide richer acoustic diversity necessary for robust Bayesian distribution estimation, we augment our training data by synthesizing multiple audio samples for each textual input using six publicly available pretrained TTS models, namely CosyVoice2 (Du et al., 2024) (LLM + flow matching-based streaming TTS), IndexTTS (Deng et al., 2025) (AR discrete acoustic token TTS), SparkTTS (Wang et al., 2025c) (LLM-based single-stage TTS), F5TTS (Chen et al., 2024a) (flow matching-based efficient TTS), MaskGCT (Wang et al., 2024) (masked generative NAR TTS), and XTTS-v2 (Casanova et al., 2024) (cross-lingual expressive TTS). Including the TTS-synthesized data, the total dataset amounts to 4,817 hours of speech. For consistency, all synthesized audio samples generated by these models are also resampled to 16 kHz.
C.2 Model Configurations
@@
C.3 Training Details of Main Results
We train BELLE, BELLE-stream and reproduce MELLE on the data in Sec.5.1, following the training strategy of combining various data sources within each batch (see Sec. 4.4), with predefined training weights that the original Librispeech audio is weighted by 0.22 and each synthesized audio source (six TTS models) is weighted by 0.13, which ensures the original Librispeech audio receives approximately twice the weighting of synthesized audio, reflecting relative considerations of synthetic audio quality. Models are trained by AdamW optimiser with a total batch size of about 160K frames distributed across 16 NVIDIA A800 GPUs. BELLE and MELLE’s training proceeds for 450K updates, where the learning rate is first linearly warmed up to a peak value of $5\times 10^{-4}$ over the initial 10% of training steps and thereafter linearly decayed to zero. Regarding the hyperparameters for BELLE, we set $\lambda=0.5$ in Eqn. equation 10 and $\lambda_{\text{samp}}=0.2$, $\lambda_{\text{flux}}=0.5$ in Eqn. equation 8. As for MELLE, our hyperparameter settings strictly follow the original MELLE paper, where $\lambda_{\text{samp}}=0.1$, $\lambda_{\text{flux}}=0.5$.
BELLE-stream is initialized from the trained BELLE model, trained with a batch size of 80K frames across 8 NVIDIA A800 GPUs for 150K updates, using $\lambda_{\text{flux}}=0.1$ while keeping all other hyperparameters identical to BELLE. In our implementation, we set $S_{\mathrm{text}}=20$ and $S_{\mathrm{audio}}=50$, corresponding to an audio chunk duration of approximately $0.8$ seconds. Here, $S_{\mathrm{text}}$ is computed from the number of phonemes obtained after G2P conversion, while $S_{\mathrm{audio}}$ is determined from the number of mel-spectrogram frames. Let $L_{\mathrm{text}}$ and $L_{\mathrm{audio}}$ denote the total phoneme count and the total mel-frame count of an utterance, respectively. During training, we filter out audio samples whose ratio $L_{\mathrm{audio}}:L_{\mathrm{text}}$ is less than $2.5$, ensuring that each text chunk retains sufficient corresponding audio frames for effective streaming generation. We observed that the synthesis quality deteriorates with speech prompt. To address this, we first investigate the potential of BELLE for streaming synthesis in a single-speaker scenario, finetuning BELLE-stream using recordings from a single speaker of Librispeech to fix its voice timbre.
C.4 Training Details of Analysis Results
@@ C.5 Evaluation Settings
For subjective evaluation, we obtain MOS and SMOS scores via a crowdsourcing platform. MOS for evaluating overall speech quality, and SMOS for measuring speaker similarity between the generated audio and the prompt. From the audio samples generated under the Cross-sentence setting, we select one sample per speaker, resulting in a total of 40 audio samples. All samples from different systems are resampled to 16kHz to ensure a fair comparison. We evaluate MOS and SMOS following the detailed procedure described in Appendix  F. Note that BELLE-stream is fine-tuned to a fixed speaker timbre and does not use an audio prompt; therefore, it is not evaluated for SMOS or SIM metrics, and its results are reported only under the Cross-sentence setting.
@@
Appendix D Details about Bayesian Evidential Learning
@@
Appendix E MELLE
E.1 Gaussian Sampling in MELLE
@@ E.2 Sampling Loss
Appendix F Subjective Evaluation

### v2

@@ D.1 Training Data
To provide richer acoustic diversity necessary for robust Bayesian distribution estimation, we augment our training data by synthesizing multiple audio samples for each textual input using six publicly available pretrained TTS models, namely CosyVoice2 (Du et al., 2024) (LLM + flow matching-based streaming TTS), IndexTTS (Deng et al., 2025) (AR discrete acoustic token TTS), SparkTTS (Wang et al., 2025c) (LLM-based single-stage TTS), F5TTS (Chen et al., 2024) (flow matching-based efficient TTS), MaskGCT (Wang et al., 2024) (masked generative NAR TTS), and XTTS-v2 (Casanova et al., 2024) (cross-lingual expressive TTS). Including the TTS-synthesized data, the total dataset amounts to 4,817 hours of speech. For consistency, all synthesized audio samples generated by these models are also resampled to 16 kHz.
D.2 Model Configurations
@@
D.3 Training Details of Main Results
We train BELLE, BELLE-stream and reproduce MELLE on the data in Sec.5.1, following the training strategy of combining various data sources within each batch (see Sec. 4.4), with predefined training weights that the original Librispeech audio is weighted by 0.22 and each synthesized audio source (six TTS models) is weighted by 0.13, which ensures the original Librispeech audio receives approximately twice the weighting of synthesized audio, reflecting relative considerations of synthetic audio quality. Models are trained by AdamW optimiser with a total batch size of about 160K frames distributed across 16 NVIDIA A800 GPUs. BELLE and MELLE’s training proceeds for 450K updates, where the learning rate is first linearly warmed up to a peak value of $5\times 10^{-4}$ over the initial 10% of training steps and thereafter linearly decayed to zero. Regarding the hyperparameters for BELLE, we set $\lambda=0.5$ in Eqn. (13) and $\lambda_{\text{samp}}=0.2$, $\lambda_{\text{flux}}=0.5$ in Eqn. (11). As for MELLE, our hyperparameter settings strictly follow the original MELLE paper, where $\lambda_{\text{samp}}=0.1$, $\lambda_{\text{flux}}=0.5$.
BELLE-stream is initialized from the trained BELLE model, trained with a batch size of 80K frames across 8 NVIDIA A800 GPUs for 150K updates, using $\lambda_{\text{flux}}=0.1$ while keeping all other hyperparameters identical to BELLE. In our implementation, we set $S_{\mathrm{text}}=20$ and $S_{\mathrm{audio}}=50$, corresponding to an audio chunk duration of approximately $0.8$ seconds. Here, $S_{\mathrm{text}}$ is computed from the number of phonemes obtained after G2P conversion, while $S_{\mathrm{audio}}$ is determined from the number of mel-spectrogram frames. Let $L_{\mathrm{text}}$ and $L_{\mathrm{audio}}$ denote the total phoneme count and the total mel-frame count of an utterance, respectively. During training, we filter out audio samples whose ratio $L_{\mathrm{audio}}:L_{\mathrm{text}}$ is less than $2.5$, ensuring that each text chunk retains sufficient corresponding audio frames for effective streaming generation.
D.4 Training Details of Analysis Results
@@ D.5 Evaluation Settings
For subjective evaluation, we obtain MOS and SMOS scores via a crowdsourcing platform. MOS for evaluating overall speech quality, and SMOS for measuring speaker similarity between the generated audio and the prompt. From the audio samples generated under the Cross-sentence setting, we select one sample per speaker, resulting in a total of 40 audio samples. All samples from different systems are resampled to 16kHz to ensure a fair comparison. We evaluate MOS and SMOS following the detailed procedure described in Appendix  H. Note that BELLE-stream is fine-tuned to a fixed speaker timbre and does not use an audio prompt; therefore, it is not evaluated for SMOS or SIM metrics, and its results are reported only under the Cross-sentence setting.
@@
Appendix E Details about Bayesian Evidential Learning
@@
Appendix F MELLE
F.1 Gaussian Sampling in MELLE
@@ F.2 Sampling Loss
Appendix G Additional Experimental Analysis
In this section, we provide further analysis regarding the hyperparameter choices in the multi-teacher framework and the sampling strategy.
G.1 Weight Allocation in Multi-Teacher Learning
In our multi-teacher learning framework, we assign specific weights to the losses derived from different teacher models. The allocation criteria reflect our prioritization of the original ground-truth audio. Specifically, we aim to ensure that the original LibriSpeech audio receives approximately twice the weight of any individual synthetic audio source. Given 6 synthetic sources, our final configuration assigns a weight of 0.22 to LibriSpeech and 0.13 to each synthetic source (totaling $\approx 1.0$).
To validate this choice, we conducted an ablation study on a reduced-scale dataset (consistent with Sec. 6.3). As shown in Table 6, the configuration with weights 0.22/0.13 yields the optimal balance, particularly minimizing WER. Deviating from this ratio (either increasing or decreasing the weight of real data) leads to a degradation in performance.
Table 6: Ablation study on teacher weight allocation. “Weight (Real)” refers to the original LibriSpeech data, and “Weight (Syn)” refers to each of the 6 synthetic sources.
Weight (Real) Weight (Syn) Continuation Cross-Sentence
WER-H $\downarrow$ SIM-o $\uparrow$ WER-H $\downarrow$ SIM-o $\uparrow$
0.16 0.14 2.64 0.415 5.17 0.453
0.22 0.13 2.57 0.409 4.73 0.449
0.28 0.12 2.78 0.406 5.47 0.444
0.34 0.11 2.76 0.405 5.47 0.446
G.2 Trade-off Between Diversity and Robustness
In Table 3 of the main text, we demonstrated that doubling the sampling parameter $\beta$ (denoted as BELLE ($\beta*2$)) significantly enhances the diversity of the generated speech (higher Cosine distance, L1, and L2 scores). However, in generative modeling, there is typically a trade-off between diversity and robustness/accuracy. To quantify this, we evaluated the WER and SIM metrics for the $\beta*2$ setting.
As shown in Table 7, while scaling $\beta$ improves diversity, it results in a moderate increase in WER (from 2.45% to 3.33%) and a slight decrease in speaker similarity. This indicates that a larger $\beta$ encourages the model to explore the “tails” of the predicted distribution, producing more varied and expressive prosody at the cost of slightly reduced stability. This trade-off is expected and allows users to tune $\beta$ according to their specific application requirements (e.g., favoring stability for reading tasks vs. diversity for creative tasks).
Table 7: Impact of doubling the sampling parameter $\beta$ on performance metrics (Cross-Sentence setting). This should be viewed in conjunction with Table 3 which shows the corresponding increase in diversity.
System WER-C $\downarrow$ SIM-o $\uparrow$
BELLE (Default) 2.45 0.641
BELLE ($\beta*2$) 3.33 0.612
G.3 The effects of the sampling module and flux loss
Table 8: Ablation study examining the effects of the sampling module and flux loss on BELLE GT + 2 TTS. ✓ indicates enabled, ✗ indicates disabled.
Sampling Flux Loss Continuation Cross-Sentence
WER-C WER-H SIM-r SIM-o WER-C WER-H SIM-r SIM-o
4.30 5.00 0.382 0.344 12.04 13.07 0.392 0.351
2.50 3.25 0.388 0.347 5.61 6.37 0.417 0.372
2.04 2.66 0.398 0.362 4.12 4.84 0.433 0.395
An ablation study is conducted to investigate how the sampling module and flux loss affect the performance of the BELLE model. Due to limited computational resources, the analysis experiments are also conducted on the smaller-scale dataset. Our experiments are based on the BELLE GT + 2 TTS setting. As demonstrated in Table 8, the sampling module plays a more crucial role than flux loss. Removing the flux loss results in only a slight performance degradation, whereas removing the sampling module leads to a severe performance drop. This indicates that Bayesian sampling plays a crucial role in the effectiveness of BELLE.
Appendix H Subjective Evaluation